OpenStudy (anonymous):

Let A, B and C be the angles of a triangle , Prove that tan (B+C) + tan (C+A) + tan (A+B) / tan(π-A) + tan(2π-B) + tan(3π-C) = 1

hartnn (hartnn):

Do u know what is the sum of all angles of the triangle?? A+B+C=?

OpenStudy (anonymous):

no

hartnn (hartnn):

ok,so i will tell you that sum of all the angles of a triangle is 180 degrees or pi radians that is A+B+C=pi from this can you find what is (pi-A) ?

OpenStudy (anonymous):

cot A ?

OpenStudy (anonymous):

 so, A+B+C=pi Thus, pi-A = B+C